OpenStudy (anonymous):

A large punch bowl holds 3.35 kg of lemonade (which is essentially water) at 20.0° C. A 1.62-kg ice cube at −10.2° C is placed in the lemonade. What is the final temperature of the system, and the amount of ice (if any) remaining? Ignore any heat exchange with the bowl or the surroundings. Answer must be in: ______degrees C ______ kg For kg, the answer is not zero Please please show all work ( even the algebra) =)

OpenStudy (anonymous):

energy will be xchanged M1C(T3-T1) = M2C(T2-T1)

OpenStudy (anonymous):

I keep trying to use that setup and I keep getting the wrong answers :(

OpenStudy (anonymous):

well C gets cancelled 3.35 (20-T) = 1.623 (T+10.2)

OpenStudy (anonymous):

I solved for T and got 10.14 and tried to enter it to my hw and it said it was wrong :/

OpenStudy (anonymous):

well no idea then, sorry

OpenStudy (anonymous):

Thank you for trying

OpenStudy (experimentx):

add latent heat of fusion

OpenStudy (experimentx):

and google for sp. heat capacity of ice

OpenStudy (anonymous):

I found that latent heat of fusion for water/lemonade is 33.5 x 10^4, Specific heat of water is 4186, and sp. heat for ice is 2090...but I don't know where to put them in the equation

OpenStudy (experimentx):

what's the answer 0 C??

OpenStudy (anonymous):

that's why I am having so much trouble with it bc there no example like it, not even in the book, I haven't tried 0 yet i will right now but I have no idea what to input for kg

OpenStudy (anonymous):

so I will try putting 0 degree C, but do you know how I can figure out the mass of ice ?

OpenStudy (anonymous):

It's correct! Thank you so so so much!!!!

OpenStudy (experimentx):

i believe you understood what i did . you are welcome

OpenStudy (anonymous):

yes I do, I will rework it again just in case! Thank you again =)

OpenStudy (experimentx):

just think what will happen ... (first the ice will come at 0 C) ... still there is some heat remaining in the water ... that heat will melt ice.

OpenStudy (anonymous):

I think I got more caught up in the math than the concept, thank you for clarifying everything!